# Energy decay of nonlocal viscoelastic equations with nonlinear damping and polynomial nonlinearity

### v1

@@ 1. Introduction
{$&u_{tt}-\mathrm{div}(A(x)\nabla u)+\int_{0}^{t}f(t-s)\,\mathrm{div}(a(x)\nabla u(s))\,\mathrm{d}s\\ &\quad+b(x)h(u_{t})=k(x)|u|^{p-2}u\end{aligned}&\quad&\mbox{in }\Omega\times\mathbb{R}_{+},\\ &u=u^{0},\ u_{t}=u^{1}&\quad&\mbox{in }\Omega\times\{0\},\\ &u=0&\quad&\mbox{on }\partial\Omega\times\mathbb{R}_{+},\end{aligned}\right.$ (1.1)
@@
[…]
@@
Over the past few decades, significant results have been established regarding well-posedness and energy decays of nonlinear wave equations with either weak or strong damping. As a pioneering work, Webb [19] first investigated the well-posedness of solutions through the operator theory and examined the long-term behavior of these solutions using Lyapunov stability techniques for the semilinear wave equation with strong damping and a forcing term. Gazzola and Squassina [8] studied the semilinear wave equation with strong frictional damping and established both existence and non-existence results for solutions when the initial energy does not exceed the depth of the potential well. Further, they discovered the blow-up phenomena by potential well methods along with Levine’s concavity techniques. Subsequently, Ma and Fang [15] demonstrated both existence and non-existence of global weak solutions and derived energy decay estimates by combining potential well methods with a logarithmic Sobolev inequality applicable to the viscoelastic wave equations. Lian and Xu [13] further explored global existence, energy decay and infinite-time blow-up using methodologies similarly to [15]. Di, Shang and Song [6] discussed the semilinear wave equation with logarithmic nonlinearity and strong damping. They proved global well-posedness along with polynomial or exponential stability through potential methods combined with Lyapunov skills, and also considered various blow-up scenarios. Cui and Chai [5] investigated the energy decay for variable coefficient wave equations with logarithmic nonlinearity by applying vector assumptions and multiplicative methods. Moreover, Yang and Zhong [20] investigated the global well-posedness and blow-up phenomena for a strongly damped wave equation with a time-varying source and singular dissipation. They established both local well-posedness and globally existing solutions based on cut-off techniques, multiplier methods, contraction mapping principles and modified potential approaches. Meanwhile, they also derived the blow-up results of solutions with arbitrary positive initial energy and the lifespan of the blow-up solutions are derived.
@@
The remainder of this paper is organized as follows. In Section 2, we specify assumptions on coefficients and prepare prelimineary lemmas including some elementary estimates of energies the global well-posedness of solutions. In Section 3, we state the two main results of this article, namely, the polynomial energy decay in Theorem 3.1 as well as a refined decay estimate in Theorem 3.2 under a weak convexity assumption on the kernel function $f(t)$. Further, we construct several examples to illustrate that Theorem 3.2 does imply faster decay rates in some cases. Subsequently, Sections 45 are devoted to the proofs of Theorems 3.13.2, respectively.
@@ 2. Preliminary and well-posedness
[…]
@@ 2.1. Some assumptions and definitions
(A3) The function $h\in C(\mathbb{R})$ is non-decreasing and satisfies
@@
(A4) The function $f\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is non-increasing and satisfies
@@
Thanks to assumptions (A2)–(A4), we see that $E^{\prime}(t)\leq 0$, that is, the energy $E(t)$ is non-increasing.
2.2. Some lemmas and well-posedness
@@ 3. Main results and examples
Here $\xi\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is non-increasing and $G\in C^{2}([0,f(0)];\overline{\mathbb{R}_{+}})$ is a strictly increasing and strictly convex function satisfying $G(0)=G^{\prime}(0)=0$.
@@
Let $q=1$ and $f(t)=\alpha\exp(-t^{\beta})$, where $\beta\in(0,1)$ and $a>0$ are again small constants so that $f$ satisfies (A4). Then this time $f$ satisfies (A5) with $\xi(t)=1$ and
@@
we see that $G$ also achieves the requirement in (A5) on $[0,\alpha]$. Then we deal with $G_{1}$ as
@@
For $I_{1}(t)$, we separate $u(s)=u(t)+(u(s)-u(t))$ and use assumptions (A2) and (A4) to estimate
@@
×$\qquad\>\>\times\widetilde{a}\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x+I_{9}(t)+I_{10}(t)$
@@
For $I_{5}(t)$, we separate $u(s)=u(t)+(u(s)-u(t))$ and use assumptions (A2) and (A4) to estimate
@@
where we used Lemma 2.1 with $r=2(p-1)$ and recall $K=\|k\|_{\infty}$. For the first term in (4.18), we combine (2.12) with (2.8) to deduce
@@
Now introduce an auxiliary function $\varphi:\overline{\mathbb{R}_{+}}\longrightarrow\overline{\mathbb{R}_{+}}$ which is non-increasing and locally absolutely continuous. For the constant $t_{0}>0$ in Case 1 and any $\tau>t_{0}$, we multiply (4.24) by $\varphi(t)$ and integrate over $(t_{0},\tau)$ to deduce
@@
Taking advantage of (2.8), (2.12) and (3.2) in Theorem 3.1, we invoke the non-negativity and the monotonicity of $E(t)$ to dominate
@@
turns out to be non-increasing. Therefore, we arrive at

### v2

@@ 1. Introduction
$\left\{\begin{aligned} &\begin{aligned} &u_{tt}-\mathrm{div}(A(x)\nabla u)+\int_{0}^{t}f(t-s)\,\mathrm{div}(a(x)\nabla u(s))\,\mathrm{d}s\\ &\quad+b(x)h(u_{t})=k(x)|u|^{p-2}u\end{aligned}&\quad&\mbox{in }\Omega\times\mathbb{R}_{+},\\ &u=u^{0},\ u_{t}=u^{1}&\quad&\mbox{in }\Omega\times\{0\},\\ &u=0&\quad&\mbox{on }\partial\Omega\times\mathbb{R}_{+},\end{aligned}\right$. (1.1)
@@
$\left\{\begin{aligned} &p>2,&\quad&n=1,2,\\ &2<p\leq\frac{2n-2}{n-2},&\quad&n\geq 3.\end{aligned}\right$.
@@
Over the past few decades, significant results have been established regarding well-posedness and energy decays of nonlinear wave equations with either weak or strong damping. As a pioneering work, Webb [19] first investigated the well-posedness of solutions through the operator theory and examined the long-term behavior of these solutions using Lyapunov stability techniques for the semilinear wave equation with strong damping and a forcing term. Gazzola and Squassina [8] studied the semilinear wave equation with strong frictional damping and established both existence and nonexistence results for solutions when the initial energy does not exceed the depth of the potential well. Further, they discovered the blow-up phenomena by potential well methods along with Levine’s concavity techniques. Subsequently, Ma and Fang [15] demonstrated both existence and nonexistence of global weak solutions and derived energy decay estimates by combining potential well methods with a logarithmic Sobolev inequality applicable to the viscoelastic wave equations. Lian and Xu [13] further explored global existence, energy decay and infinite-time blow-up using methodologies similarly to [15]. Di, Shang and Song [6] discussed the semilinear wave equation with logarithmic nonlinearity and strong damping. They proved global well-posedness along with polynomial or exponential stability through potential methods combined with Lyapunov skills, and also considered various blow-up scenarios. Cui and Chai [5] investigated the energy decay for variable coefficient wave equations with logarithmic nonlinearity by applying vector assumptions and multiplicative methods. Moreover, Yang and Zhong [20] investigated the global well-posedness and blow-up phenomena for a strongly damped wave equation with a time-varying source and singular dissipation. They established both local well-posedness and globally existing solutions based on cut-off techniques, multiplier methods, contraction mapping principles and modified potential approaches. Meanwhile, they also derived the blow-up results of solutions with arbitrary positive initial energy and the lifespan of the blow-up solutions are derived.
@@
The remainder of this paper is organized as follows. In Section 2, we specify assumptions on coefficients and prepare preliminary lemmas including some elementary estimates of energies the global well-posedness of solutions. In Section 3, we state the two main results of this article, namely, the polynomial energy decay in Theorem 3.1 as well as a refined decay estimate in Theorem 3.2 under a weak convexity assumption on the kernel function $f(t)$. Further, we construct several examples to illustrate that Theorem 3.2 does imply faster decay rates in some cases. Subsequently, Sections 45 are devoted to the proofs of Theorems 3.13.2, respectively.
@@ 2. Preliminary and well-posedness
Throughout this article, we denote the norm of the Lebesgue space $L^{p}(\Omega)$ ($1\leq p\leq\infty$) by $\|\cdot\|_{p}$, and the inner product of $L^{2}(\Omega)$ by (,)($\cdot\,$$\cdot\,$).
@@ 2.1. Some assumptions and definitions
(A3) The function $h\in C(\mathbb{R})$ is nondecreasing and satisfies
@@
(A4) The function $f\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is nonincreasing and satisfies
@@
Thanks to assumptions (A2)–(A4), we see that $E^{\prime}(t)\leq 0$, that is, the energy $E(t)$ is nonincreasing.
$2.$ Some lemmas and well-posedness
@@ 3. Main results and examples
Here $\xi\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is nonincreasing and $G\in C^{2}([0,f(0)];\overline{\mathbb{R}_{+}})$ is a strictly increasing and strictly convex function satisfying $G(0)=G^{\prime}(0)=0$.
@@
Let $q=1$ and $f(t)=\alpha\exp(-t^{\beta})$, where $\beta\in(0,1)$ and $\alpha>0$ are again small constants so that $f$ satisfies (A4). Then this time $f$ satisfies (A5) with $\xi(t)=1$ and
@@
we see that $G$ also achieves the requirement in (A5) on $0,\alpha]$. Then we deal with $G_{1}$ as
@@
For $I_{1}(t)$, we separate u(s)=u(t)+(u(s)u(t))u(s)=u(t)+(u(s)-u(t)) and use assumptions (A2) and (A4) to estimate
@@
$\qquad\>\>\times\widetilde{a}\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x+I_{9}(t)+I_{10}(t)$
@@
For $I_{5}(t)$, we separate u(s)=u(t)+(u(s)u(t))u(s)=u(t)+(u(s)-u(t)) and use assumptions (A2) and (A4) to estimate
@@
where we used Lemma 2.1 with r=2(p1)r=2(p-1) and recall $K=\|k\|_{\infty}$. For the first term in (4.18), we combine (2.12) with (2.8) to deduce
@@
Now introduce an auxiliary function $\varphi:\overline{\mathbb{R}_{+}}\longrightarrow\overline{\mathbb{R}_{+}}$ which is nonincreasing and locally absolutely continuous. For the constant $t_{0}>0$ in Case 1 and any $\tau>t_{0}$, we multiply (4.24) by $\varphi(t)$ and integrate over $(t_{0},\tau)$ to deduce
@@
Taking advantage of (2.8), (2.12) and (3.2) in Theorem 3.1, we invoke the nonnegativity and the monotonicity of $E(t)$ to dominate
@@
turns out to be nonincreasing. Therefore, we arrive at